# FedMeld: A Model-dispersal Federated Learning Framework for Space-ground Integrated Networks

### v2

@@ V-C Learning Performance
V-D Effects of Satellite Number
@@
V-E Effects of Latency Constraint
@@
V-F Effect of Data Heterogeneity
@@ Appendix A Proof of Theorem 1
By setting $\beta=\frac{5}{\mu}$ and $\gamma=\max\left\{\frac{8L}{\mu},E\right\}-1$, then $\eta_{t}$ satisfies $\eta_{t}\leq 2\eta_{t+E}$ for any step $t$. Then, we have
@@
$\leq\frac{{25B}}{{{4\mu^{2}}}}+\left({\gamma+1}\right){\Delta_{1}}.$
@@
$\begin{split}&\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{t}}}\right)}\right]-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\frac{L}{2}{\Delta_{t}}\\ &\leq\frac{L}{2}\left[\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right){\frac{{v}}{{t+\gamma}}+{2Q_{t-1}}}\right]\\ &\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({t+\gamma}\right)}}\left[{\frac{{25B}}{{{8\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}{\Delta_{1}}}\right]+L{Q_{t-1}}.\end{split}$

### v3

@@ V-D Learning Performance
V-E Effects of Satellite Number
@@
V-F Effects of Latency Constraint
@@
V-G Effect of Data Heterogeneity
@@ Appendix A Proof of Theorem 1
By setting $\beta=\frac{2}{\mu}$ and $\gamma=\max\left\{\frac{8L}{\mu},E\right\}-1$, then $\eta_{t}$ satisfies $\eta_{t}\leq 2\eta_{t+E}$ for any step $t$. Then, we have
@@
$\leq\frac{{4B}}{{{\mu^{2}}}}+\left({\gamma+1}\right){\Delta_{1}}.$
@@
$\begin{split}&\mathbb{E}\left[{F\left({{{\overline{\mathbf{w}}}_{t}}}\right)}\right]-F\left({{{\mathbf{w}}^{\star}}}\right)\leq\frac{L}{2}{\Delta_{t}}\\ &\leq\frac{L}{2}\left[\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right){\frac{{v}}{{t+\gamma}}+{2Q_{t-1}}}\right]\\ &\leq\left({\frac{1}{{1-\alpha}}-\alpha+\frac{1}{2}}\right)\frac{{L}}{{\mu\left({t+\gamma}\right)}}\left[{\frac{{2B}}{{{\mu}}}+\frac{{\mu\left({\gamma+1}\right)}}{2}{\Delta_{1}}}\right]+L{Q_{t-1}}.\end{split}$